Write the equilibrium expression (K) for each of t
Answer #1

a.

N2 + O2 = NO

balance

N2 + O2 = 2NO

K = [NO]^2 / ([N2][O2])

b.

N2O4 --> NO2

the balance

N2O4 --> 2NO2

K = [NO2]^2 / [N2O4]

c.

PBr3 + Cl2 = PCl3 + Br2

balance

2PBr3 + 3Cl2 = 2PCl3 + 3Br2

K = [PCl3]^2 [ Br2]^3 / ([PBr3]^2[Cl2]^3)